OpenStudy (anonymous):

Solve for x: (1.5)x-6=10

OpenStudy (anonymous):

Isolate x 1.5x-6=10 1.5x=16 x=16/1.5 x=10.66 repeating x=10.667 wherever you wish to stop on the decimal place.